OpenStudy (anonymous):

miss clark bought 3 cups of yogurt and 5 oranges for $2.79 then she bought 4 yogurts and 7 oranges for $3.81. find the cost of one yogurt and the cost of one orange

OpenStudy (a1234):

3y + 5o = 2.79 4y + 7o = 3.81 I'll use elimination: 12y + 20o = 11.16 12y + 21o = 11.43 -o = -.27 or o = 0.27 3y + 5(.27) = 2.79 3y + 1.35 = 2.79 3y = 1.44 y = 0.48 (0.27, 0.48) Check to see if that solution works.

OpenStudy (anonymous):

thanks

OpenStudy (a1234):

You're welcome.